# supplement

### v1

@@ C.6 Lower bound
The lower bound corresponding to Theorem 1 is obtained by the same steps as the upper bound, replacing “$\sup$” with “$\inf$”. In Cases 1 and 2, this leads to a flipped sign for the solutions of the free components in Eqs. C.3 and C.13, affecting the feasibility checks. Overall, the optimal lower bound is also obtained by case distinction:
$\underline{f}(x_{N+1})=\begin{cases}-\sqrt{K^{f}_{N+1,N+1}}\Gamma_{f},&\text{in Cases~{}1 and 4, }\\ \Phi_{N+1}\theta^{\mu}-\|P\Phi_{N+1}^{\top}\|_{P^{-1}}\sqrt{\Gamma_{w}^{2}-y^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}y+\|\theta^{\mu}\|_{P^{-1}}^{2}}&\text{in Cases 2 and 4, }\\ f^{\mu}_{\sigma^{\star}}(x_{N+1})-\sqrt{\Gamma_{f}^{2}+\frac{\Gamma_{w}^{2}}{(\sigma^{\star})^{2}}-\|g^{\mu}_{\sigma^{\star}}\|_{\mathcal{H}_{k^{f}+k^{\sigma^{\star}}}}^{2}}\sqrt{\Sigma^{f}_{\sigma^{\star}}(x_{N+1})},&\text{in Cases 3 and 4.}\end{cases}$
=$=\begin{cases}\lim_{\sigma\rightarrow\infty}\underline{f}^{\sigma}(x_{N+1}),&\text{if \mbox{$\left\|y+K^{f}_{1:N,N+1}\frac{\Gamma_{f}}{\sqrt{K^{f}_{N+1,N+1}}}\right\|_{(K^{w}_{1:N,1:N})^{-1}}^{2}\leq\Gamma_{w}^{2}$} },\\ \lim_{\sigma\rightarrow 0}\underline{f}^{\sigma}(x_{N+1}),&\text{if \mbox{$\left\|\theta^{\mu}-\frac{P\Phi_{N+1}^{\top}}{\|P\Phi_{N+1}^{\top}\|_{P^{-1}}}\sqrt{\Gamma_{w}^{2}-y^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}y+\|\theta^{\mu}\|_{P^{-1}}^{2}}\right\|_{2}^{2}\leq\Gamma_{f}^{2}$}, }\\ \sup_{\sigma\in(0,\infty)}\underline{f}^{\sigma}(x_{N+1}),&\text{otherwise.}\end{cases}$
Analogous to Section C.5, by replacing “$\inf$” with “$\sup$” and flipping the corresponding inequalities, the optimal solution is shown to be given as

### v3

@@ C.6 Lower bound
The lower bound corresponding to Theorem˜1 is obtained by the same steps as the upper bound, replacing $\sup$” with $\inf$”. In Cases 1 and 2, this leads to a flipped sign for the solutions of the free components in Eqs.˜C.3 and C.13, affecting the feasibility checks. Overall, the optimal lower bound is also obtained by case distinction:
[…]
$=\begin{cases}\lim_{\sigma\rightarrow\infty}\underline{f}^{\sigma}(x_{N+1}),&\text{if \mbox{$\left\|y+K^{f}_{1:N,N+1}\frac{\Gamma_{f}}{\sqrt{K^{f}_{N+1,N+1}}}\right\|_{(K^{w}_{1:N,1:N})^{-1}}^{2}\leq\Gamma_{w}^{2}$} },\\ \lim_{\sigma\rightarrow 0}\underline{f}^{\sigma}(x_{N+1}),&\text{if \mbox{$\left\|\theta^{\mu}-\frac{P\Phi_{N+1}^{\top}}{\|P\Phi_{N+1}^{\top}\|_{P^{-1}}}\sqrt{\Gamma_{w}^{2}-y^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}y+\|\theta^{\mu}\|_{P^{-1}}^{2}}\right\|_{2}^{2}\leq\Gamma_{f}^{2}$}, }\\ \sup_{\sigma\in(0,\infty)}\underline{f}^{\sigma}(x_{N+1}),&\text{otherwise.}\end{cases}$
Analogous to Section˜C.5, by replacing $\inf$” with $\sup$” and flipping the corresponding inequalities, the optimal solution is shown to be given as
@@
Appendix D Numerical example on safe control for uncertain nonlinear systems: Implementation details
In the following, we describe the setup of the optimization problem solved in the numerical example “Safe control for uncertain nonlinear systems” in Section˜4.3. Inserting the system dynamics (18) into the safety constraint $x(k+1)\geq(1-\gamma)x(k)$, the condition $f^{\mathrm{known}}(x(k),u(k))+f^{\mathrm{tr}}(x(k),u(k))\geq(1-\gamma)x(k)$ can be enforced robustly by utilizing the lower uncertainty bound $f^{\mu}_{\sigma}(x,u)-\beta_{\sigma}\sqrt{\Sigma_{\sigma}(x,u)}\leq f^{\mathrm{tr}}(x,u)$, leading to the tightened constraint
$(1-\gamma)x(k)$ $\leq f^{\mathrm{known}}(x(k),u(k))+f^{\mu}_{\sigma}(x(k),u(k))-\beta_{\sigma}\sqrt{\Sigma_{\sigma}(x(k),u(k))}.$
Minimization of the user-defined cost $c(x(k),u(k))$ subject to the above constraint can be achieved by solving the following optimization problem:
$\min_{x_{0},x_{1},u_{0},\sigma,s}\quad$ $c(x_{0},u_{0})+\omega s$
$\mathrm{s.t.}\quad$ $x_{0}=x(k),$
$x_{1}=f^{\mathrm{known}}(x_{0},u_{0})+f^{\mu}_{\sigma}(x_{0},u_{0}),$
$x_{1}\geq(1-\gamma)x_{0}+\beta_{\sigma}\sqrt{\Sigma_{\sigma}(x_{0},u_{0})}-s,$
$s\geq 0,$
$u_{\max}\geq u_{0}\geq u_{\min}.$
Therein, the added slack variable $s$ allows the optimizer to converge even in case the decrease condition is impossible to satisfy given the bounds $u_{\min},u_{\max}$ on the control input. A large linear penalty $\omega>0$ thereby incentivizes $s=0$, i.e., constraint satisfaction; a solution is classified as feasible if the optimal slack value $s^{\star}$ satisfies $s^{\star}\leq 10^{-6}$. The optimization problems are implemented in CasADi [Andersson et al., 2019] and solved using the interior-point optimizer IPOPT [Wächter and Biegler, 2006] on an Intel i9-7940X CPU. Table˜1 provides all parameters and expressions used for the implementation.
Parameter Value
$f^{\mathrm{known}}(x,u)$ 0.5x+u10.5x+u-1
$c(x,u)$ $(f^{\mathrm{known}}(x,u)+f^{\mu}_{\sigma}(x,u))^{2}+u^{2}$
$u_{\min}$ 2-2
$u_{\max}$ 22
$\gamma$ $0.95$
$\omega$ $10^{4}$
Table 1: Parameters and expressions for the CBF example
Additional implementation details can be found in the published source code at https://gitlab.ethz.ch/ics/bounded-energy-rkhs-bounds and at https://doi.org/10.3929/ethz-c-000785083.